OpenStudy (anonymous):

Please help, last question. Dont understand it divide 19x^4+15x^3/3x^3 Simplify your answer as much as possible.

jimthompson5910 (jim_thompson5910):

It is this right? \[\Large \frac{19x^4 + 15x^3}{3x^3}\]

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

medal plzz

jimthompson5910 (jim_thompson5910):

what you can do is break up the fraction like this \[\Large \frac{19x^4 + 15x^3}{3x^3} = \frac{19x^4}{3x^3}+\frac{15x^3}{3x^3}\] very similar to how \(\Large \frac{1}{7} + \frac{3}{7} = \frac{1+3}{7}\)

jimthompson5910 (jim_thompson5910):

then you simplify each piece

jimthompson5910 (jim_thompson5910):

or you can treat it like this \[\Large \frac{19x^4 + 15x^3}{3x^3} = \frac{x^3(19x+15)}{3x^3}\] notice how I factored out an x^3 term

OpenStudy (anonymous):

27x^4-18x^3/ 3x^2 , I got the one above wrong and skipped me to the next question. But thats the problem I am on now.

jimthompson5910 (jim_thompson5910):

what is the GCF of the numerator?